OpenStudy (anonymous):

what is the solution of this equation? 7 + 1/3x =9

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

\[7+\frac{1}{3x}=9\] \[\frac{1}{3x}=9-7\] \[\frac{1}{3x}=2\] \[1=2*3x\] \[1=6x\] \[\frac{1}{6}=x\]